OpenStudy (anonymous):

Which ordered pair is a solution of the equation y=x-4 A. (2,6) B. (6,2) C. (-2,6) D. (-6,2)

OpenStudy (anonymous):

For every coordinate you have (x,y) you can plug in these values into your equation to see if they are true. y=x-4 A. 6 = 2-4, so 6= -2 not true B. 2 = 6-4, so 2 = 2 true <------------ C. 6 = -2-4, -6=-6 not true D. 2= -6-4, 2=-10 not true

OpenStudy (anonymous):

@nikkimessina do you get it??